ft status and plans
Download
Skip this Video
Download Presentation
FT Status and Plans

Loading in 2 Seconds...

play fullscreen
1 / 4

FT Status and Plans - PowerPoint PPT Presentation


  • 107 Views
  • Uploaded on

FT Status and Plans. Overall CTS Worked to get CTS back on line Slots A4-7 have problems Had to work on SBC Don L. is continuing tests with new VRBC and VRB firmware at CTS L1 Deadtime – Extra Data in readout Put 470KOhm 0402 resistors on un-used channels 8 per MCM, 64 per board

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' FT Status and Plans' - zuri


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
ft status and plans
FT Status and Plans
  • Overall
    • CTS
      • Worked to get CTS back on line
        • Slots A4-7 have problems
        • Had to work on SBC
      • Don L. is continuing tests with new VRBC and VRB firmware at CTS
    • L1 Deadtime – Extra Data in readout
      • Put 470KOhm 0402 resistors on un-used channels
      • 8 per MCM, 64 per board
      • Verify << in process now
        • Board ready for platform
      • Set up modification line
  • Access AFE board swap
    • CFT
      • Replaced 9A5 >> dip at X2
      • Replaced 7A4
        • Inductors to look at double peds
    • CPS
      • Re-programmed CPSS personality
    • FPS
    • CTT

6 Mayl 2003

july shutdown
July Shutdown
  • General items being considered
    • Cryostat work
      • About 1 degK left in setting
      • Have been moving about 0.1
      • Will not warm up “unless” have further problems
    • Cassette work
      • Only if we warm up
      • Perhaps install ‘testing’ cassette
    • Will do
      • Unused channel fix for AFE brds
      • Check/calibrate/fix AFE lvps monitoring
      • Add SEQ crate monitoring
      • Finish LED pulser boards and cabling
      • New firmware
      • Fix slot problems – AFE & SEQ
        • 4B6 SVX Downloads
        • 6B2/3 SEQ signals @ SEQ BP
    • May do
      • AC remote shutoff for AFE lvps
      • 53MHz readout
      • Re-cable back of SEQ crate PC04
    • Will not do
  • CPS
  • FPS
  • CTT
  • FPD

Starts ~July 28 >> 4 to 6 weeks

6 Mayl 2003

july shutdown 2
July Shutdown 2
  • Details
    • How long to do?
    • Who can work on?
    • When can it be done?

Starts ~July 28 >> 4 to 6 weeks

6 Mayl 2003

merge shifts
Merge Shifts
  • Management has scheduled a merger of CFT and SMT shifts
    • Would start after shutdown
      • after MUON and CAL merge
        • Should be merged by end June
  • Can the shifts be merged
    • Reduce the amount of work the CFT shifter does
    • Increase effectiveness per minute of attention
    • Improve and condense the documentation
  • How are shifts merged
    • Need people from both groups with cross knowledge
    • Need a small core of people to get it done

6 Mayl 2003

ad